Operations Report – January 31st

Crew 220 Operations Report 31JAN2020

Sol: 12

Name of person filing report: Matt Wise

Non-nominal systems: Primary Generator, Water Heater, Hab lower deck 1 fluorescent light ballast.

Generator: Primary is inoperable. Rental was run-the following information is for the rental.

Hours run: 11.5
From what time last night: 20:00

To what time this morning: 07:30

List of any additional daytime hours when the generator was run: N/A

Solar- SOC% (Before the generator is run at night): 79%

Diesel Reading – Just over 1/4 mark

Station Propane Reading- 76%

Ethanol Free Gasoline- N/A

Water (loft tank) (gal) – 35

Water Meter (units)- 0148627,9

Water (static tank) (gal) – 43

Static to loft Pump used – Yes 38 gal

Water in GreenHab (gal): 100.04

Water in Science Dome (gal): 0

Toilet Tank Emptied – Yes

Deimos rover used: Still in the workshop


Beginning Charge:

End Charge:

Currently Charging:

Sojourner rover used: No


Beginning Charge:

End Charge:

Currently Charging:

Spirit Rover used – Yes

Starting Hours: 128.6

Beginning Charge: 100%

Ending Hours: 128.7

Ending Charge: 91%

Opportunity Rover used: Still in workshop


Beginning Charge:

Ending Charge:

Currently Charging:

Curiosity Rover used: Yes

Starting Hours: 134.6

Beginning Charge: 100%

Ending Hours: 134.7

Ending Charge: 91%

Notes on rovers: Opportunity and Deimos off-site for maintenance. Should be operational within a few sols

ATVs Used: (Honda, 300, 350.1, 350.2, 350.3): No

Reason for use:

Oil Added? Checked and added as necessary

# hours the ATV’s were used today: 0

Notes on ATVs:

Habcar used and why, where? Yes. For water resupply

CrewCar used and why, where? Yes. In SLC picking up supplies

General Notes and Comments: A small white bowl was accidentally broken while washing the evening dishes.

Summary of the internet: Nominal

Summary of Suit and Radios: 3 radios malfunctioning
Suits Nominal

Summary of Hab: Main fluorescent lights in Hab lower deck malfunctioning. The problem seems to stem from irregular electrical current flow from temp generator. Center light bar has been fixed and seems to function properly at all times, light bar near back airlock has been repaired and seems to be functioning properly, while light bar near stairs has no bulbs.

Summary of Science Dome operations: Nominal

Summary of Ram operations: Crew member using RAM for assembly of 3D printer

Summary of any observatory issues: Nominal

Summary of Health and Safety issues: MDRS crew all healthy

Questions, concerns, and requests to Mission Support: Please remove trash from rear airlock. Final trash pickup Feb 2. Resupply of toilet paper.

Copyright © The Mars Society. All rights reserved. | Main Site